Identify the center and radius of x2 + (y - 1)2 = 36
Answer by nerdybill(6948)

The standard equation for a circle is:
(x - h)^2 + (y - k)^2 = r2
where h and k are the x- and y-coordinates of the center of the circle and r is the radius.
.
Now, let's look at your problem.
x^2 + (y - 1)^2 = 36
we can rewrite it as:
(x-0)^2 + (y - 1)^2 = 6^2
.
Based on the above, we now have:
h=0
k=1
r=6
.
center = (h,k) = (0,1)
radius = r = 6